A rectangle is drawn so the width is 71 inches longer than the height. If the rectangle's diagonal measurement is 85 inches, find the height.
Give your answer rounded to 1 decimal place.

Answer by TimothyLamb(4379)

w = h + 71
d = 85 = sqrt( ww + hh )
---
85 = sqrt( ww + hh )
85^2 = ww + hh
85^2 = (h + 71)(h + 71) + hh
85^2 = hh + 142h + 71^2 + hh
2h^2 + 142h + 71^2 - 85^2 = 0
2h^2 + 142h - 2184 = 0
---
the above quadratic equation is in standard form, with a=2, b=142, and c=-2184
---
to solve the quadratic equation, plug this:
2 142 -2184
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots of the quadratic are:
h = 13
h = -84
---
the negative root doesn't make sense for a length, so use the positive root
---
Answer:
h = 13 in
w = 84 in
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php